# Influence of initial states on memory effects: A study of early-time superradiance

### v1

@@ V Conclusions and Discussions
In this paper, we propose a method to evaluate the memory effects in a quantum process conditioned on a particular system initial state. The method is based on the physical interpretations of the inequality $T(t_{2},t_{0})\neq T(t_{2},t_{1})T(t_{1},t_{0})$ [30]. Some features of the non-Markovianity measure in [30] are inherited. For example, nonzero memory effects can be characterized even in regimes where monotonic behaviors do not occur, or have not occurred yet. Besides, $0\leqslant N_{M}[\rho_{S}(t_{0})]\leqslant 1$ is satisfied without additional normalizations. This allows us to compare the memory effects of quantum systems with different dimensions (e.g., different numbers of atoms in our model). Using our method, we calculate the influence of several types of initial states on the early-time memory effects. For the (dephased, mixed) Dicke states, we obtained simple analytic solutions for the normalized memory effects in terms of $J$ and $M$, or say, $N_{e}$ and $N_{g}$. The environmental photon number as well as the degree of superradiance are also calculated. The aims of comparing the memory effects with the superradiance characters is to reveal the role of memory effects in generating superradiance, and conversely, understand the physical sources of the memory effects.
The main conclusions are as follows: For the (dephased) Dicke states, the value of memory effects is half the cavity photon number in the early-time regime. For the factorized identical states, the single-atom coherence is necessary for superradiance. Such coherence can always enhance the memory effects for $N\geqslant 1$, meanwhile, it increases the cavity photon number and the degree of superradiance for $N\geqslant 2$. For the dephased Dicke states, the coherence (entanglement of atoms) represented by $\lambda$ is also necessary for superrdiance. Meanwhile, the memory effects, the cavity photon number and the degree of superradiance increase linearly with $\lambda$ for superradiant states. In the above cases, the transitions from non-superradiant states to superradiant ones are always accompanied by the enhancement of memory effects, implying the memory effects are vital for creating superradiance. On the other hand, we find that the (change of) environmental photon number is an important (but may not be the only) source of the memory effects in the superradiance process, especially for the Dicke states.
VI ACKNOWLEDGMENTS

### v2

@@ VI Discussions and Conclusions
In this paper, we propose a method to evaluate the memory effects in a quantum process conditioned on a particular system initial state. The method is based on the physical interpretations of the inequality $T(t_{2},t_{0})\neq T(t_{2},t_{1})T(t_{1},t_{0})$ [35]. Some features of the non-Markovianity measure in Ref. [35] are inherited. For example, nonzero memory effects can be characterized even in regimes where nonmonotonic behaviors do not occur, or have not occurred yet. Besides, $0\leqslant N_{M}[\rho_{S}(t_{0})]\leqslant 1$ is satisfied without additional normalizations. This allows us to compare the memory effects of quantum systems with different dimensions (e.g., different numbers of atoms in our model). Using our method, we calculate the influence of two types of initial states on the memory effects as well as the superradiance in different regimes. The characteristics of the memory effects and the superradiance are compared to reveal the role of memory effects in superradiance, and conversely, to understand the physical sources of the memory effects.
In the early-time (Zeno) regime and the (early-stage) near-Markovian regime where the initial states do not change so much, the main observations and conclusions are as follows. The memory effects for the (dephased) Dicke states are fully determined by the cavity photon number (connected to the atom radiation intensity) and they are proportional to $(gt)^{2}(J+M)[(J-M)\lambda+1]$ for quadratic dynamics, or $(g/\gamma)^{2}(J+M)[(J-M)\lambda+1]$ in the Markovian limit. Meanwhile, the degree of superradiance is equal to the degree memory-effect-enhancement (from independent radiation to collective radiation) given by $(J-M)\lambda+1$. For the factorized identical states, the memory effects, the radiation intensity and the degree of superradiance can all be enhanced by the single-atom coherence. The memory effects are closely related to the cavity photon number, meanwhile, the superradiance ($S>1$) of an initial state is accompanied by the enhancement of memory effects (by a common cavity). These results demonstrate that the memory effects are closely related to the radiation intensity. Correspondingly, the enhancement of memory effects by a common environment plays an important role on the establishment of superradiance, especially for the Dicke states. On the other hand, these results imply that the (change of) cavity photon number is one fundamental source of memory effects, especially for the Dicke states. Besides, the entanglement in the Dicke states and the single-atom coherence in the factorized identical state are necessary for the existence of superradiance and important for the memory-effect-enhancement.
For the whole dynamics in a strongly non-Markovian regime, an initial state may evolves dramatically during the radiation lifetime, making its influence weaker. However, simulations demonstrate that the relationships between the radiation intensity (degree of superradiance) and the memory effects (memory-effect-enhancement) are basically consistent with those in the other two regimes. The results in different regimes demonstrate that the characteristics of memory effects and superradiance in long-time dynamics can be reflected by its early-time behaviours. In addition, our work also provides other ways to measure the memory effects through the atom excitation number or the cavity photon number.
VII ACKNOWLEDGMENTS